# AutoWebGLM: Bootstrap And Reinforce A Large Language Model-based Web Navigating Agent

### v1

@@ Abstract
Large language models (LLMs) have fueled many intelligent agent tasks, such as web navigation—but most existing agents perform far from satisfying in real-world webpages due to three factors: (1) the versatility of actions on webpages, (2) HTML text exceeding model processing capacity, and (3) the complexity of decision-making due to the open-domain nature of web. In light of the challenge, we develop AutoWebGLM, a GPT-4-outperforming automated web navigation agent built upon ChatGLM3-6B. Inspired by human browsing patterns, we design an HTML simplification algorithm to represent webpages, preserving vital information succinctly. We employ a hybrid human-AI method to build web browsing data for curriculum training. Then, we bootstrap the model by reinforcement learning and rejection sampling to further facilitate webpage comprehension, browser operations, and efficient task decomposition by itself. For testing, we establish a bilingual benchmark—AutoWebBench—for real-world web browsing tasks. We evaluate AutoWebGLM across diverse web navigation benchmarks, revealing its improvements but also underlying challenges to tackle real environments. Related code, model, and data will be released at https://github.com/THUDM/AutoWebGLM.
11footnotetext: HL, XL and ILI make equal contribution. Emails: {laihy23,rongyl20}@mails.tsinghua.edu.cn, shawliu9@gmail.com22footnotetext: Work done when HL, ILI, SY, YC, PS, HY, and HCZ interned at Zhipu AI.
Figure 1: AutoWebGLM on various web browsing benchmarks. Despite improvements, there is still much gap between it and human performance on challenging real-world missions.
Figure 2: Examples of AutoWebGLM’s execution on four example user tasks.
1 Introduction
@@
The concept of autonomous digital agents as helpful assistants is an enticing prospect. Enhanced by LLMs’ formidable comprehension and response capabilities [1; 35; 36; 45; 46; 34], we can envision various scenarios unimaginable before. For instance, an LLM-based agent could support a daily routine that summarizes the online news for us during breakfast. This integration of LLMs into everyday tasks heralds a significant shift in how we interact with technology, optimizing our efficiency and redefining the boundaries of machine-assisted productivity [41; 37; 21].
@@
  • Lack of Unified Action Space: A universal and convenient action space covering all necessary task executions on browser across various websites is absent.
@@
  • Lack of Webpage Simplification Method: The diversity and complexity of webpages and their tendentious verbosity pose a significant challenge for LLMs to comprehend the content and carry out correct operations. Token length of content-rich webpages can usually reach 30k and over.
@@
  • Lack of High-quality Training Trajectories: There are limited high-quality trajectories for training a strong LLM-based web agent. Existing trained agents notably lack the capability for correct inference and self-checking on web tasks. Once caught in an erroneous loop, they struggle to rectify the issue promptly.
@@
In this work, we introduce AutoWebGLM, a deployable webpage browsing agent based on the open ChatGLM3-6B model [45]. Different from its predecessor WebGLM [20] that focuses on retrieval augmented web-scale question answering, AutoWebGLM is dedicated to autonomously accomplish complex real-world missions via navigating and operating on real web browsers like humans. We propose various efficient data strategies to support the swift construction of a sizeable, reliable training dataset while state-of-the-art models cannot reliably complete data annotation tasks [47]. Furthermore, by leveraging supervised [30] and reinforcement learning methods [32], we train AutoWebGLM on top of the collected web agent dataset to achieve performance superiority on general webpage browsing tasks. A step further, we employ rejection sampling finetuning (RFT)  [36] for lifelong learning in specific web environments, enabling the agent to become an expert in a particular domain.
@@
We develop a Chrome extension based on AutoWebGLM (See Figure 2 for examples). Throughout our experiments, it can reason and perform operations on various websites to complete user tasks accurately, making it practically applicable to real-world services. In addition, we construct the first bilingual (English and Chinese) webpage browsing evaluation dataset, given that websites from different regions have substantial stylistic variations.
In conclusion, we make the following contributions in this paper:
@@
  • We perform experiments to demonstrate that AutoWebGLM with 6B parameters achieves performance comparable to the most advanced LLM-based agents, and more importantly, it reaches a practically usable level for real-world web tasks.
2 Method
@@
Figure 3: The System Architecture of AutoWebGLM. Our system comprises two key components: browsing framework and LM agent. The browsing framework (left) uses various web processing modules to organize concise HTML and other information for the LM agent to make decisions that are then executed by an automated browsing program. The LM agent (right) learns from data procured from diverse sources. It further employs RL and RFT to bootstrap itself, thus enhancing web browsing capabilities.
2.1 Problem Setup
@@
We consider web browsing tasks as a sequence decision-making process. The state, denoted as $S$, includes the current page status (such as HTML, URL, and Window Position). The action set $A$ contains all potential browsing operations, including click, type, scroll, etc. See complete operations in Table 1.
@@
The state’s transition is determined by the webpage’s current state and the agent’s output action. The task will end when the agent outputs finish or reaches the maximum number of interactions.
$S_{t+1}=T(S_{t},A_{t})$
During the decision-making process, the function $\phi$ updates the historical information based on the previous history $H_{t-1}$, the most recent action $A_{t-1}$, and the current state $S_{t}$.
@@
$(S_{t+1},H_{t+1})=(T(S_{t},A_{t}),\phi(H_{t},A_{t},S_{t+1})),A_{t}=\pi(S_{t}\mid H_{t})$
2.2 The AutoWebGLM Framework
A comprehensive, precise observation and action space is vital for constructing a robust web browsing framework. These spaces standardize the conversion of varied data sources into a uniform format.
2.2.1 Observation space
HTML. The HTML webpages are vast and complex, so it is necessary to simplify them before inputting them into the model. The simplification process aims to extract essential information while eliminating redundant or disruptive elements that could hinder the model’s understanding. Throughout this process, the HTML’s basic structure and significant content information must be retained to enable the model to comprehend and utilize this information for effective web browsing. The algorithm in Appendix A can efficiently convert a tree of elements into a concise representation. We can use the processing techniques to streamline the original HTML format into a more understandable structure for the model to interpret and manage, improving model effectiveness in web browsing tasks.
2.2.2 Action space
As the approach of this work is to build a language model-based web browsing agent, we focus on operational possibilities when constructing the action space. On an extensive summary of experiences in the real task execution process, we define a complete and self-consistent action space (in Table 1) formulated as function calls [21; 11] for the language model to act in the web browsing world. We design our prompt input in Section D.
Table 1: Action space for AutoWebGLM to interact through.
@@ 2.2.2 Action space
2.3 Data Preparation
As illustrated in Figure 4, we suggest a hybrid human-AI Data Construction method to create our training data to deal with these challenges. After careful consideration, we categorize our data into two types for construction:
2.3.1 Web Recognition & Simple Task Operation Construction
Figure 4: Data Construction. Data construction is divided into two main stages; the first stage is webpage recognition tasks and simple tasks operation construction, and the second stage is complex tasks construction.
2.3.2 Complex Task Operation Construction
Figure 5: Dataset proportion of splits within our training data.
We first designed 50 complex tasks for each website using the prompting technique referring to Evol-Instruct [42], from which about 20 feasible tasks were manually selected and labeled. For operation sequence, due to the high complexity of the tasks, even the most advanced LLMs cannot complete the task with satisfactory accuracy. Therefore, we leveraged manual annotations to capture web task executions via a browser plugin that records actions during website tasks. Chain-of-thought [40] reasoning has been proven to improve task comprehension and model performance [17; 39] significantly. However, leveraging human annotators to document their intent and reasoning during web browsing is inefficient. To improve the CoT construction process, we used GPT-4 as the operational intent predictor. Our first approach of iterative step-by-step creation proved to generate weak operational links and incurred high API costs due to data construction. To address this, we employed a global thought chain prompting method, where all operations and critical HTML segments are inputted into a trace. Then, we prompted GPT-4 to output intentions for each step. This method improves the accuracy and cohesion of each step, thus forming highly relevant, consistent thought chains.
2.3.3 AutoWebBench Construction
Figure 6: The Training Procedure. First, the model learns webpage interpretation and operation via curriculum learning. Next, it self-samples training data, learning from its mistakes. Finally, it self-plays in the environment, becoming a domain expert.
2.4 Training
We train the model through three steps illustrated in Figure 6.
2.4.1 Step 1: Curriculum Learning
$\mathcal{L_{\rm SFT}}(\pi_{\theta})=-\mathbb{E}_{(x,y)\thicksim\mathcal{D}}\left[\log\pi_{\theta}(y\mid x)\right]$ (1)
This approach enhances the model’s comprehension of webpages and its capability as an agent to perform operations within the environments. Significantly, we use curriculum learning (CL), which mimics the human learning process, advocating for models to start learning from easy samples and gradually advance to complex ones. It has been demonstrated in prior works[6; 38] to improve model capabilities substantially.
2.4.2 Step 2: Reinforcement Learning
After constructing contrastive data $D_{\text{Const.}}$, we employ the DPO[32] training approach to make $M_{\text{SFT}}$ learn from its mistakes and further enhance its capabilities. During the training, we found that the direct use of DPO loss led to instability. To mitigate this issue, we propose including SFT loss to stabilize the reinforcement learning process and increase the number of training steps while ensuring no loss of the original model’s natural language and agent abilities, achieving a more robust model $M_{\text{DPO}}$:
$\begin{array}[]{l}\mathcal{L}_{\rm DPO}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\thicksim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right]\end{array}$ (2)
$\mathcal{L}_{\rm SFT}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\thicksim\mathcal{D}}\left[\log\pi_{\theta}(y_{w}\mid x)\right]$ (3)
$\mathcal{L}_{\rm Total}=\lambda\cdot\mathcal{L}_{\rm DPO}+\mathcal{L}_{\rm SFT}$ (4)
2.4.3 Step 3: Rejection Sampling Finetuning
3 Experiments
Table 2: Step Success Rates of different models on AutoWebBench. All models are tested with in-context learning prompts presented in Appendix E.
Model #Params English Chinese
Cross-Task Cross-Domain Cross-Task Cross-Domain
GPT-3.5-Turbo [29] N/A 12.1 6.4 13.5 10.8
GPT-4 [1] N/A 38.6 39.7 36.7 36.3
Claude2 [3] N/A 13.2 8.1 13.0 7.9
LLaMA2 [36] 7B 3.3 2.5 - -
LLaMA2 [36] 70B 8.3 8.9 - -
Qwen [4] 7B 9.0 7.6 9.1 7.5
AutoWebGLM 6B 64.8 58.6 65.4 61.8
Table 3: Step Success Rates on Mind2Web. † indicates that top-10 candidates were used for this test, otherwise top-50 was used. * indicates model’s individual finetuning on corresponding train set.
Model #Params Modality Cross-Task Cross-Website Cross-Domain Avg.
GPT-3.5-Turbo [29] N/A Text 17.4 16.2 18.6 17.4
GPT-4 [1] N/A Text 36.2 30.1 26.4 30.9
Flan-T5-XL [22] 3B Text 52.0 38.9 39.6 43.5
Html-T5-XL [12] 540+3B Text 71.5 62.2 67.1 66.9
LLaMA2 [36] 7B Text 52.7 47.1 50.3 50.1
LLaMA2 [36] 70B Text 55.8 51.6 55.7 54.4
Qwen-VL [5] 9.6B Image & Text 12.6 10.1 8.0 10.2
SeeClick [8] 9.6B Image & Text 23.7 18.8 20.2 20.9
CogAgent [14] 18B Image & Text 62.3 54.0 59.4 58.2
AutoWebGLM 6B Text 66.4 56.4 55.8 59.5
3.1 Main Results
Beyond AutoWebBench, we also test AutoWebGLM over three other established web navigating benchmarks: Mind2Web [9], MiniWoB++ [19], and WebArena [47].
AutoWebBench. As discussed in Section 2.3.3, We divide the test set into four splits: Chinese, English, in-domain, and out-of-domain, for evaluation purposes. We use the Step Success Rate (SSR) as our evaluation metric. All models are evaluated with an in-context learning prompt as described in Appendix E. The results are in Table 2. As discerned from the table, AutoWebGLM, after multi-task training, excels in predicting general user operation patterns, aligning well with user operations. In contrast, other baselines, in the absence of sufficient training, struggle to accurately learn user operations across different real-world websites based on webpage content and task descriptions.
Mind2Web [9]. We use the settings from Mind2Web with SSR as our primary evaluation metric. To compare the model fairly, we utilize the MindAct framework provided by Mind2Web to evaluate the model’s performance. The results are in Table 3. We obtained the baseline results from references [4; 9; 12; 14; 8].
Figure 7: Results on MiniWoB++ and WebArena. * indicates model’s individual finetuning on task-related datasets.
Model Size MiniWoB++ WebArena
GPT-3.5-Turbo [29] N/A 13.4 6.2
GPT-4 [1] N/A 32.1 14.4
Text-Bison-001 [2] N/A - 5.1
LLaMA2 [36] 7B 42.8 1.2
LLaMA2 [36] 70B 47.1 0.6
WebN-T5-XL [13] 3B 48.4 -
Html-T5-XL [12] 543B 85.6 -
Lemur [43] 70B - 5.3
AutoWebGLM 6B 89.3 18.2
MiniWoB++ [19] & WebArena [47]. For MiniWob++, following the experimental setup from WebAgent [12], we test MiniWoB++ with 56 tasks by running 100 evaluation episodes per task to evaluate model capabilities. For WebArena, we integrate our HTML parser module and action execution module into the WebArena environment to make it compatible with our system. The results are in Table 7. For the WebArena baselines, the results are derived from the references [47; 43; 44]. Regarding the MinWob++ baselines, some of the results come from the references [12]. LLaMA2 results are obtained through training and evaluation on the MinWob++ dataset.
3.2 Ablation Study
Table 4: Ablation study. AutoWebBench and WebArena do not have a training set, while the RFT stage is only suitable for sampling in the environment, so we represent them by "-".
Method AutoWebBench Mind2Web MiniWob++ WebArena
Training Data Ablation
Only Train Set - 48.1 44.3 -
+) Stage1 23.5 48.4 48.3 2.5
+) Stage2 60.2 55.2 78.9 7.6
+) Stage1+2 61.8 56.7 81.7 8.3
Training Strategy Ablation
SFT 61.8 56.7 81.7 8.3
+) DPO 62.7 59.5 80.8 8.5
+) RFT - - 89.3 18.2
AutoWebGLM 62.7 59.5 89.3 18.2
3.3 Case Study and Error Analysis
To assess the effectiveness of our model, we conduct a series of case studies covering a range of web-based tasks, including everyday use, leisure and relaxation, and academic research, covering the typical range of web requirements. Our system achieves satisfactory results in most scenarios, with several specific cases detailed in the appendix G.
Table 5: Error Distribution in Web Task Automation
Error Type Proportion
Hallucinations 44%
Poor Graphical Recognition 28%
Misinterpretation of Task Context 20%
Pop-Up Interruption 8%
4 Related Work
Language Models (LMs). Large language models (LLMs), such as GPT-4 [1], Claude-2 [3], LLaMA-2 [35], GLM-130B [45; 10], OPT [46], and BLOOM [34], have accumulated extensive knowledge in various natural language processing tasks. However, due to the high cost of deploying such large language models, smaller models with lower costs and comparable capabilities are usually preferred. Many open-source projects, such as LLaMA2-7B [35] and ChatGLM3-6B [45], have demonstrated strong performance to large language models in some domains.
Benchmarks for Web Navigation. The primary web browsing evaluation datasets provide a variety of evaluation metrics. MiniWoB++ [16] provides several simulated web environments, with tasks primarily to evaluate the model’s ability to interact with webpage components. However, with the increasing demand for complex web operation capabilities, Mind2Web [9] and WebArena [47] have been created. Mind2Web is an offline evaluation set for complex web browsing that provides several metrics. The evaluation method is straightforward and commonly used for model evaluations. In contrast, the WebArena benchmark, based on real websites, creates multiple virtual environments and uses various evaluation methods to assess the task completion rate, making it more suitable for real-world task completion evaluation.
Agents for Web Automation. Previous work such as WebGPT [27] and WebGLM [20] combined LLMs with web environments. However, their primary application was question-answering (QA) tasks [33; 28; 7; 18], utilizing internet resources to answer user queries. Recent works [25; 14; 8; 43] focus more on executing complex operations or interactive tasks. Specifically, MindAct [9] works by filtering webpage elements and selecting the element through multiple rounds of multiple-choice questions. It often requires more than ten model calls to complete a single web operation, which could be more efficient. On the other hand, WebAgent [12] uses HTML-T5 to process the observation space’s content, including HTML, previous operations, and user instructions. It uses the Flan-U-Plam model to generate code to control webpages, exhibiting excellent web browsing performance. However, it faces deployment challenges due to the size of the Flan-U-Plam model, which is 540B scale. AutoWebGLM, based solely on a single ChatGLM3-6B, has a robust web browsing capability comparable to WebAgent, demonstrating high value for practical deployment.
Prompt-based Data Construction Methods. Constructing data through prompts has recently gained significant traction [39; 15; 31; 26]. This approach leverages language models to generate synthetic data for training. A notable example is Evol-Instruct [42; 23], inspired by the theory of evolution, demonstrating the effectiveness of using LLMs to generate diverse and complex instructions for various tasks. Additionally, some researchers explore the potential of generating data in a zero-shot setting, where the model produces data for tasks it has yet to be explicitly trained on [24], highlighting the versatility of prompt-based data construction. These methodologies rapidly evolve, offering a promising avenue for data generation in various domains, especially where traditional data collection methods could be more practical and sufficient.
5 Conclusion
In this work, we present AutoWebGLM, an advanced language model-based agent exhibiting robust performance in various autonomous web navigation benchmarks. Our model addresses extant LLM limitations and simplifies webpages by effectively controlling HTML text length and handling the web’s open-domain nature. We strategically employ curriculum learning, reinforcement learning, and rejection sampling finetuning to enhance webpage comprehension and browser operation learning. We also introduce a unique bilingual web browsing benchmark— that lays a solid foundation for future research. Our findings represent significant progress in utilizing LLMs for intelligent agents.
Appendix A HTML Prunning Pseudo Code

### v2

# AutoWebGLM: A Large Language Model-based Web Navigating Agent

@@ Abstract.
Large language models (LLMs) have fueled many intelligent web agents, but most existing ones perform far from satisfying in real-world web navigation tasks due to three factors: (1) the complexity of HTML text data (2) versatility of actions on webpages, and (3) task difficulty due to the open-domain nature of the web. In light of these challenges, we develop the open AutoWebGLM based on ChatGLM3-6B. AutoWebGLM can serve as a powerful automated web navigation agent that outperform GPT-4. Inspired by human browsing patterns, we first design an HTML simplification algorithm to represent webpages with vital information preserved succinctly. We then employ a hybrid human-AI method to build web browsing data for curriculum training. Finally, we bootstrap the model by reinforcement learning and rejection sampling to further facilitate webpage comprehension, browser operations, and efficient task decomposition by itself. For comprehensive evaluation, we establish a bilingual benchmark—AutoWebBench—for real-world web navigation tasks. We evaluate AutoWebGLM across diverse web navigation benchmarks, demonstrating its potential to tackle challenging tasks in real environments. Related code, model, and data are released at https://github.com/THUDM/AutoWebGLM.
ChatGLM, Large Language Model, LLM Agent, Web Agent, Reinforcement Learning, Rejection Sampling Finetuning
copyright: rightsretainedjournalyear: 2024doi: 10.1145/3637528.3671620conference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Intelligent agents
Figure 1. The performance of AutoWebGLM on various web browsing tasks in comparison with GPT-4 and open LLMs.
Figure 2. Examples of AutoWebGLM’s execution on four user tasks.
1. Introduction
@@
The concept of autonomous digital agents as helpful assistants is an enticing prospect. Enhanced by LLMs’ formidable comprehension and response capabilities (Achiam et al., 2023; Touvron et al., 2023a, b; Zeng et al., 2022; Zhang et al., 2022; Scao et al., 2022), we can envision various scenarios previously unimaginable. For instance, an LLM-based agent could support a daily routine that summarizes the online news from the open web for us. This integration of LLMs into everyday tasks heralds a significant shift in how we interact with machines, optimizing our efficiency and redefining the boundaries of machine-assisted productivity (Xi et al., 2023; Wang et al., 2023a).
Tremendous efforts have been underway to construct auto web agents. One is AutoGPT, a popular open-source project that utilizes ChatGPT (OpenAI, 2022) to integrate LLMs with predetermined tools such as web and local file browsing. Meanwhile, the development of agent-centric LLMs has gained significant momentum (Yao et al., 2022; Press et al., 2022; Wang et al., 2023b; Hong et al., 2023). Nevertheless, the majority of existing web agents are to date severely restricted in terms of practical applications, predominantly due to the following challenges:
@@
  • A universal action space covering all necessary task executions across various websites is absent.
@@
  • The diversity and complexity of webpages and their tendentious verbosity pose a significant challenge for LLMs to comprehend the content and carry out correct operations accurately.
@@
  • Existing agents notably lack the capability for correct inference and self-checking on web tasks. Once caught in an erroneous loop, they struggle to rectify the issue promptly.
@@
In this work, we introduce AutoWebGLM for building webpage navigation agents. It is built upon the open ChatGLM3-6B model (Zeng et al., 2022). First, we propose various efficient data strategies to support the swift construction of a sizeable, reliable training dataset while state-of-the-art models cannot reliably complete data annotation tasks (Zhou et al., 2023). Furthermore, by leveraging supervised (Ouyang et al., 2022) and reinforcement learning methods (Rafailov et al., 2023), we train AutoWebGLM on top of the collected web agent dataset to achieve performance superiority on general webpage browsing tasks. A step further, we employ rejection sampling finetuning (RFT)  (Touvron et al., 2023b) for lifelong learning in specific web environments, enabling the agent to excel in a particular domain.
@@
We develop and deploy a Chrome extension based on AutoWebGLM (See Figure 2 for examples). Throughout our experiments, it can reason and perform operations on various websites to complete user tasks accurately, making it practically applicable to real-world services. In addition, we construct the first bilingual (English and Chinese) webpage browsing evaluation dataset to build AutoWebBench, given that websites from different regions have substantial stylistic variations. In conclusion, we make the following contributions in this paper:
@@
  • We perform experiments to demonstrate that AutoWebGLM with six billion parameters achieves performance comparable to the most advanced LLM-based agents, and more importantly, it reaches a practically usable level for real-world web tasks.
2. Related Work
Language Models (LMs). Large language models (LLMs) (Zhao et al., 2023), such as GPT-4 (Achiam et al., 2023), Claude-2 (Anthropic, 2023), LLaMA-2 (Touvron et al., 2023a), ChatGLM (Zeng et al., 2022; Du et al., 2022), OPT (Zhang et al., 2022), and BLOOM (Scao et al., 2022), have accumulated extensive knowledge in various natural language processing tasks. However, due to the high cost of deploying such large language models, smaller models with lower costs and comparable capabilities are usually preferred. Many open-source projects, such as LLaMA-2-7B (Touvron et al., 2023a) and ChatGLM3-6B (Zeng et al., 2022), have demonstrated strong performance to large language models in some domains.
Benchmarks for Web Navigation. The primary web browsing evaluation datasets provide a variety of evaluation metrics. MiniWoB++ (Humphreys et al., 2022) provides several simulated web environments, with tasks primarily to evaluate the model’s ability to interact with webpage components. However, with the increasing demand for complex web operation capabilities, Mind2Web (Deng et al., 2023) and WebArena (Zhou et al., 2023) have been created. Mind2Web is an offline evaluation set for complex web browsing that provides several metrics. The evaluation method is straightforward and commonly used for model evaluations. In contrast, the WebArena benchmark, based on real websites, creates multiple virtual environments and uses various evaluation methods to assess the task completion rate, making it more suitable for real-world task completion evaluation.
Agents for Web Automation. Previous work such as WebGPT (Nakano et al., 2021) and WebGLM (Liu et al., 2023) combined LLMs with web environments. However, their primary application was question-answering (QA) tasks (Rajpurkar et al., 2016; Nguyen et al., 2016; Berant et al., 2013; Kwiatkowski et al., 2019), utilizing internet resources to answer user queries. Recent works (Mishra et al., 2019; Hong et al., 2023; Cheng et al., 2024; Xu et al., 2023a) focus more on executing complex operations or interactive tasks. A fundamental aspect of web browsing tasks involves a comprehensive understanding of HTML. StructGPT (Jiang et al., 2023) explores methodologies to enhance the zero-shot reasoning ability of LLMs in handling structured data. Specifically, MindAct (Deng et al., 2023) works by filtering webpage elements and selecting the element through multiple rounds of multiple-choice questions. It often requires more than ten model calls to complete a single web operation, which could be more efficient. On the other hand, WebAgent (Gur et al., 2023) uses HTML-T5 to process the observation space’s content, including HTML, previous operations, and user instructions. It uses the Flan-U-Plam model to generate code to control webpages, exhibiting excellent web browsing performance. However, it faces deployment challenges due to the size of the Flan-U-Plam model, which is 540B scale. AutoWebGLM, based solely on a single ChatGLM3-6B, has a robust web browsing capability comparable to WebAgent, demonstrating high value for practical deployment.
Prompt-based Data Construction Methods. Constructing data through prompts has recently gained significant traction (Wang et al., 2022; Honovich et al., 2022; Peng et al., 2023; Mukherjee et al., 2023; Cheng et al., 2023). This approach leverages language models to generate synthetic data for training. A notable example is Evol-Instruct (Xu et al., 2023b; Luo et al., 2023), inspired by the theory of evolution, demonstrating the effectiveness of using LLMs to generate diverse and complex instructions for various tasks. Additionally, some researchers explore the potential of generating data in a zero-shot setting, where the model produces data for tasks it has yet to be explicitly trained on (Meng et al., 2022), highlighting the versatility of prompt-based data construction. These methodologies rapidly evolve, offering a promising avenue for data generation in various domains, especially where traditional data collection methods could be more practical and sufficient..
Rejection Sampling Finetuning. The methodology of Rejection Sampling Finetuning (RFT) (Yuan et al., 2023) employs a supervised learning model to generate and collect accurate reasoning paths, subsequently utilized as an augmented finetuning dataset. Using RFT to expand the dataset with diverse reasoning paths can boost the mathematical performance of LLMs. Our experiments show that RFT can also be effectively implemented in web page browsing tasks, significantly increasing professional capabilities within specific environments.
@@ 2. Related Work
Figure 3. The System Architecture of AutoWebGLM. Our system comprises two key components: interaction framework and LM agent. The LM agent learns from data procured from diverse sources. It further employs RL and RFT to bootstrap itself, thus enhancing web browsing capabilities. The interaction framework uses various web processing modules to organize concise HTML and other information for the LM agent to make decisions that are then executed by an automated browsing program.
3. AutoWebGLM as A Web Agent
3.1. Problem Setup
@@
We consider web browsing tasks as a sequential decision-making process. The state, denoted as $S$, includes the current page status, such as HTML, URL, and Window Position. The action set $A$ contains all potential browsing operations, including click, type, scroll, etc (See complete operations in Table 1).
@@
The state’s transition is determined by the webpage’s current state and the agent’s output action. During the decision-making process, the function $\phi$ updates the historical information based on the previous history $H_{t-1}$, the most recent action $A_{t-1}$, and the current state $S_{t}$.
@@
$(S_{t+1},H_{t+1})=(T(S_{t},A_{t}),\phi(H_{t},A_{t},S_{t+1}))$
$A_{t}=\pi(S_{t}\mid H_{t})$
$S_{t+1}=T(S_{t},A_{t})$
Table 1. Action space for AutoWebGLM to interact through.
@@
3.2. The AutoWebGLM Framework
A comprehensive, precise observation and action space is vital for constructing a robust web browsing framework. These spaces standardize the conversion of varied data sources into a uniform format. We discuss our designs in the following:
3.2.1. Observation space